# Semantic Web: Past, Present, and Future

### v1

@@ 6.1 Reasoning over Graph Data
OWL offers a variety of other constructs for the description of classes, relationships, and concrete facts. For example, OWL allows the declaration of transitive relations and inverse relations. For example, the relation “is-member” is inverse to “has-member”. OWL reasoning allows, among other things, consistency checking of an ontology or checking the satisfiability of classes [87]. A class is satisfiable if there can be instances of that class.
For a detailed discussion about OWL reasoning, we refer to the literature such as [87, 16]. Different reasoners for OWL have seen widespread adoption in the community such as the well-known Pellet [^50] https://github.com/stardog-union/pellet and Hermit [67]. Finally, a combination of description logic and rules is also possible. For example, Motik et al. [116] presented a combination of description logic and rules that allows tractable inference on OWL ontologies.
@@ 6.2 Linking of Objects and Concepts
In the Semantic Web, it cannot be assumed that two URIs refer to two different real-world objects (cf. unique name assumption in Section 5.2). A URI by itself, or in itself, has no identity [75]. Rather, the identity or interpretation of a URI is revealed by the context in which it is used on the Semantic Web. Determining whether or not two URIs refer to the same entity is not a simple task and has been studied extensively in data mining and language understanding in the past. For example, to identify whether or not the author names of research papers refer to the same person, it is often not sufficient to resolve the name, venue, title, and co-authors [97]. The process of determining the identity of a resource is often referred to as entity resolution [97], coreference resolution [166], object identification [134], and normalization [166, 167]. Correctly determining the identity of entities on the Web is important as more and more records appear on the Web and this presents a significant hurdle for very large Semantic Web applications [66].
@@
To address this, a number of services exist that can recognize entities and determine their identity: Thomson Reuters offers OpenCalais [^51] https://www.refinitiv.com/en/products/intelligent-tagging-text-analytics, a service that can link natural language text to other resources using entity recognition. Another commercial tool that allows for extracting knowledge graphs from text is provided by DiffBot. [^52] https://www.diffbot.com/ Recently, the language model ChatGPT has been compared to the specialized entity and relation extraction tool REBEL [30] for the task of creating knowledge graphs from sustainability-related text [156]. The experiments suggest that large language models improve the accuracy of creating knowledge graphs [156]. The sameAs [^53] http://sameas.org/ service aims to detect duplicate resources on the Semantic Web using the OWL relationship owl:sameAs. This can be used to resolve coreferences between different datasets. For example, for the query with the URI http://dbpedia.org/resource/ABBA, a list of over 100 URIs is returned that also references the music group ABBA. One of them is BBC with the resource http://www.bbc.co.uk/music/artists/d87e52c5-bb8d-4da8-b941-9f4928627dc8#artist.
@@
Furthermore, the problem of schema matching [167] is very related to the problem of entity resolution, co-reference resolution, and normalization. The goal of schema matching is to address the question of how to integrate data [167], which is non-trivial even for small schemas. In the Semantic Web, schema matching means the matching of different ontologies, respectively the concepts defined in these ontologies. Various (semi-)automatic or machine learning techniques for matching ontologies have been developed in the past [54, 50, 25]. Core ontologies as illustrated in Figure 4.3 represent generic modeling frameworks for integration and alignment with other ontologies. In addition, core ontologies can also integrate Linked Open Data, which typically contains no or very little schema information. The YAGO ontology [151] was generated from the fusion of Wikipedia and Wordnet using rule-based and heuristic methods. A manual assessment showed an accuracy of 95%.
@@ 7 Querying of Linked Data
Queries over Linked Data can be processed using link traversal [79], i. e., the query processor would use one of those IRIs given directly in the query as starting point and query the respective source for more triples involving the IRI. By iteratively doing this for more IRIs and with respect to the graph pattern defined in the query, a local set of triples is collected over which the given query can be evaluated.
@@ 7.1 Basic Query Processing
As an example, let us consider the query illustrated in Figures 8 and 9 that we want to execute over our example MusicBrainz graph from Section 2. We are now interested in the musicians of ABBA who are also members of other bands. If we follow the Linked Data principles and evaluate the query using link traversal [79], this would mean first querying for triples including the IRI that represents ABBA, then navigating to the individual band members, and then following the links to all of the members’ bands and query more relevant triples.
@@
Similar to relational database systems, there exist several dedicated graph stores (aka triple stores) that are optimized for RDF graphs and evaluating SPARQL queries. Some of the most popular triple stores are RDF4J [29], Jena [169], Virtuoso [^58] http://virtuoso.openlinksw.com, and GraphDB [^59] https://graphdb.ontotext.com/. They are building upon concepts and techniques known from relational database systems [90, 117] and expand them with graph-specific optimizations [164, 149, 73, 107, 55].
@@ 7.2 Entailment Regimes and Query Processing
Although both approaches yield equivalent answers for a given SPARQL query, their performance can diverge significantly. Materialized RDF query processing may outperform on-the-fly execution of the rewritten query, but it may consume more memory [68]. Nevertheless, various optimization techniques have been proposed to mitigate the overhead caused by the on-the-fly evaluation of entailment regimes [157]. These optimizations are required particularly in the presence of the owl:sameAs. This predicate corresponds to logical equivalence and involves the application of the Leibniz Inference Rule [71] to deduce all the equivalent triples entailed by equivalent resources based on owl:sameAs relation. This process may lead to many intermediate results, impacting the query engine’s performance. Xiao et al. [171] propose query rewriting techniques to efficiently evaluate SPARQL queries with owl:sameAs employing equivalent SQL queries.
@@ 7.3 Federated Query Processing
Federations provide another perspective on querying linked data over multiple sources. A federation of knowledge graphs shares common entities while potentially providing different perspectives on those entities. Each knowledge graph within the federation operates autonomously and can be accessed through various Web interfaces, such as SPARQL endpoints or Linked Data Fragments (LDFs) [162]. SPARQL endpoints offer users the ability to execute any SPARQL query against multiple SPARQL endpoints. In contrast, LDFs enable access to specific graph patterns, such as triple patterns [161] or star-shaped graph patterns [6], allowing retrieval of fragments from an RDF knowledge graph. A star-shaped subquery is a conjunction of triple patterns in a SPARQL query that share the same subject variable [164]. An LDF client can submit requests to a server, which then delivers results based on a data shipping policy and partitions results into batches of specified page sizes. Query processing in a federation of graphs differs from querying a single source because it enables real-time data integration of graphs from multiple sources. For example, Figure 10 depicts a SPARQL query whose execution requires the evaluation of subqueries over three knowledge graphs: a Cancer Knowledge Graph (CKG) [7], DBpedia, and Wikidata. This query could not be executed over a single data source unless the three knowledge graphs were physically materialized into one. Subqueries with a specific shape (e. g., star-shaped subqueries) need to be identified and posed against the knowledge graph(s) that is able to answer a particular part of the query. The federated query engine has to decompose input queries into these subqueries, find a plan to execute them and collect and merge the answers from the subqueries to produce a federated answer.
@@
A federated SPARQL query engine typically follows a mediator and wrapper architecture, which has been established in previous research [168, 172]. Wrappers play a crucial role in translating SPARQL subqueries into requests sent to SPARQL endpoints, while also converting the endpoint responses into internal structures that the query engine can process. The mediator, on the other hand, is responsible for rewriting the original queries into subqueries that can be executed by the data sources within the federation. Additionally, the mediator collects and merges the results obtained from evaluating the subqueries to produce the final answer to the federated query. Essentially, the mediator consists of three main components:
@@
  • Source selection and query decomposition. This component decomposes queries into subqueries and selects the appropriate graphs (sources) capable of executing each subquery. Simple subqueries typically consist of a list of triple patterns that can be evaluated against at least one graph. Formally, source selection corresponds to the problem of finding the minimal number of knowledge graphs from the federation that can produce a complete answer to the input query. On the other hand, query decomposition requires partitioning the triple patterns of a query into a minimal number of subqueries, such that each subquery can be executed over at least one of the selected knowledge graphs. Commonly, federated query engines follow heuristic-based methods to solve these two problems. For example, for query decomposition, heuristics based on exclusive groups [145] or star-shaped subqueries [164, 163, 113] enable to efficiently solve source selection and query decomposition in queries free of general predicates (e. g., owl:sameAs or rdf:type).
@@
    Moreover, more general approaches (e. g., Endris et al. [51]) resort to metadata describing the star-shaped patterns existing in a knowledge graph to perform query decomposition and source selection accurately. An exclusive group corresponds to a set of triple patterns (corresponding to a conjunction of the triples) in a SPARQL query that can be exclusively executed over one data source, as in FedX [145]. Such a technique is implemented in the SPARQL engine FedX [145], which also optimizes the execution of star-shaped subqueries by identifying groups of conjunctive triple patterns that a data source can execute independently from others.
@@
  • Query optimizer. This component identifies execution plans by combining star-shaped subqueries (SSQs) and utilizing physical operators implemented by the query engine. Formally, optimizing a query corresponds to the problem of finding a physical plan for the query that minimizes the values of a utility function (e. g., execution time or memory consumption). To maximize the utility function, query optimizers consider plans with different orders of executing operators, alternative implementations of operators, such as joins, as well as particular execution alternatives for certain query types, e. g., queries involving aggregation [93]. In general, finding an optimal solution is computationally intractable [92], while the problems of constructing a bushy tree plan [143] and finding an optimal query decomposition over the graphs [163] are NP-Hard. A bushy tree plan is a query execution plan that represents a query as a tree structure with multiple branches or subqueries, which can also be bushy-tree plans. Query plans can be generated following the traditional optimize-then-execute paradigm or re-optimize and adapt a plan on the fly according to the conditions and availability of selected graphs [52]. Alternatively, the query optimizer may resort to a cost model to guide the search on the space of query plans and identify the one that minimizes the values of the utility function [113].
@@
  • Query engine. This component of a federated query engine implements the physical operators necessary to combine tuples obtained from the graphs. These physical operators are designed to support logical SPARQL operations such as JOIN, UNION, or OPTIONAL [126]. Physical operators can be empowered to adapt execution schedulers to the current conditions of a group of selected graphs. Thus, adaptivity can be achieved at the intra-operator level, where the operators can detect when graphs become blocked or data traffic bursts. Additionally, intra-operator opportunistically produce results as quickly as data arrives from the graphs, and can produce results incrementally. Some opportunistic approaches [5, 89, 61, 4] combine producing results quickly in an incremental fashion with greedy source selection so that the system stops querying additional graphs once the user’s wishes, e. g., in terms of the minimum number of obtained results, are fulfilled. The physical operators implemented by the SPARQL federated query engine ANAPSID [5] implement intra-operator adaptivity, enabling results generation even when the graphs became blocked. On the other hand, inter-operator adaptive strategies can produce an answer as soon as it is computed and can keep producing intermediate results even when data from a source becomes blocked. Acosta and Vidal [4] propose adaptive query processing techniques that enable the re-ordering of a query execution plan to adjust execution schedulers to unexpected environmental conditions (e. g., changes at rates at which tuples arrive from graphs or a graph’s availability).
@@
The problem of query execution is defined in terms of data shipping strategies that are able to distribute query load between a server-client architecture of Linked Data Fragments. State of the art approaches tackle this issue by employing graph partitioning techniques to generate graph pattern fragments. These fragments aim to improve query performance, scalability, and workload balance. SAGE [110], smart-KG [14], brTPF [80], SkyTPF [101], SPF [6], TPF [162], and WiseKG [13] are exemplary approaches for accessing graph patterns of varying shapes. These interfaces for Linked Data Fragments may significantly accelerate query execution, reduce workload, and minimize data transfer. Moreover, query rewriting based on these fragments correctly produces all the answers of the rewritten query, i. e., the methods are sound and complete. However, it is important to note that there exists a trade-off between memory consumption and execution time improvement so that storage requirements (e. g., smart-KG) can be twice as high in comparison to other approaches (e. g., SAGE, TPF, or SPF), due to the need to store partitions and their associated metadata.
Because of the variety of available interfaces, some works have proposed methods for heterogeneous federations [83, 112, 111, 38], where the goal is to find efficient query execution plans in consideration of the constraints as well as strengths of the available interface.

### v2

@@ 6.1 Reasoning over Graph Data
OWL offers a variety of other constructs for the description of classes, relationships, and concrete facts. For example, OWL allows the declaration of transitive relations and inverse relations. For example, the relation “is-member” is inverse to “has-member”. OWL reasoning allows, among other things, consistency checking of an ontology or checking the satisfiability of classes [DBLP:reference/sp/HorrocksP11]. A class is satisfiable if there can be instances of that class.
For a detailed discussion about OWL reasoning, we refer to the literature such as [DBLP:reference/sp/HorrocksP11, DBLP:journals/sLogica/BaaderS01]. Different reasoners for OWL have seen widespread adoption in the community such as the well-known Pellet [^50] https://github.com/stardog-union/pellet and Hermit [DBLP:journals/jar/GlimmHMSW14]. Finally, a combination of description logic and rules is also possible. For example, Motik et al. [DBLP:conf/semweb/MotikSS04] presented a combination of description logic and rules that allows tractable inference on OWL ontologies.
@@ 6.2 Linking of Objects and Concepts
In the Semantic Web, it cannot be assumed that two URIs refer to two different real-world objects (cf. unique name assumption in Section 5.2). A URI by itself, or in itself, has no identity [HalpinP09]. Rather, the identity or interpretation of a URI is revealed by the context in which it is used on the Semantic Web. Determining whether or not two URIs refer to the same entity is not a simple task and has been studied extensively in data mining and language understanding in the past. For example, to identify whether or not the author names of research papers refer to the same person, it is often not sufficient to resolve the name, venue, title, and co-authors [PallikaEtAlCoreference2007]. The process of determining the identity of a resource is often referred to as entity resolution [PallikaEtAlCoreference2007], coreference resolution [WickEtAlCoreference], object identification [RendleS06], and normalization [WickEtAlCoreference, WickKDD2008]. Correctly determining the identity of entities on the Web is important as more and more records appear on the Web and this presents a significant hurdle for very large Semantic Web applications [GlaserLOD2009].
@@
To address this, a number of services exist that can recognize entities and determine their identity: Thomson Reuters offers OpenCalais [^51] https://www.refinitiv.com/en/products/intelligent-tagging-text-analytics, a service that can link natural language text to other resources using entity recognition. Another commercial tool that allows for extracting knowledge graphs from text is provided by DiffBot. [^52] https://www.diffbot.com/ Recently, the language model ChatGPT has been compared to the specialized entity and relation extraction tool REBEL [DBLP:conf/emnlp/CabotN21] for the task of creating knowledge graphs from sustainability-related text [DBLP:journals/corr/abs-2305-04676]. The experiments suggest that large language models improve the accuracy of creating knowledge graphs [DBLP:journals/corr/abs-2305-04676]. The sameAs [^53] https://www.sameas.cc/ service aims to detect duplicate resources on the Semantic Web using the OWL relationship owl:sameAs. This can be used to resolve coreferences between different datasets. For example, for the query with the URI http://dbpedia.org/resource/ABBA, a list of over 100 URIs is returned that also references the music group ABBA. One of them is BBC with the resource http://www.bbc.co.uk/music/artists/d87e52c5-bb8d-4da8-b941-9f4928627dc8#artist.
@@
Furthermore, the problem of schema matching [WickKDD2008] is very related to the problem of entity resolution, co-reference resolution, and normalization. The goal of schema matching is to address the question of how to integrate data [WickKDD2008], which is non-trivial even for small schemas. In the Semantic Web, schema matching means the matching of different ontologies, respectively the concepts defined in these ontologies. Various (semi-)automatic or machine learning techniques for matching ontologies have been developed in the past [euzenat2007b, Ehrig07, Blomqvist09]. Core ontologies as illustrated in Figure 4.3 represent generic modeling frameworks for integration and alignment with other ontologies. In addition, core ontologies can also integrate Linked Open Data, which typically contains no or very little schema information. The YAGO ontology [YagoWWW07] was generated from the fusion of Wikipedia and Wordnet using rule-based and heuristic methods. A manual assessment showed an accuracy of 95%.
@@ 7 Querying of Linked Data
Queries over Linked Data can be processed using link traversal [DBLP:conf/esws/Hartig11], i. e., the query processor would use one of those IRIs given directly in the query as starting point and query the respective source for more triples involving the IRI. By iteratively doing this for more IRIs and with respect to the graph pattern defined in the query, a local set of triples is collected over which the given query can be evaluated.
@@ 7.1 Basic Query Processing
As an example, let us consider the query illustrated in Figures 8 and 9 that we want to execute over our example MusicBrainz graph from Section 2. We are now interested in the musicians of ABBA who are also members of other bands. If we follow the Linked Data principles and evaluate the query using link traversal [DBLP:conf/esws/Hartig11], this would mean first querying for triples including the IRI that represents ABBA, then navigating to the individual band members, and then following the links to all of the members’ bands and query more relevant triples.
@@
Similar to relational database systems, there exist several dedicated graph stores (aka triple stores) that are optimized for RDF graphs and evaluating SPARQL queries. Some of the most popular triple stores are RDF4J [iswc-Broekstra02sesame], Jena [DBLP:conf/semweb/WilkinsonSKR03], Virtuoso [^58] http://virtuoso.openlinksw.com, and GraphDB [^59] https://graphdb.ontotext.com/. They are building upon concepts and techniques known from relational database systems [DBLP:conf/rweb/HoseSTW11, DBLP:journals/pvldb/NeumannW08] and expand them with graph-specific optimizations [DBLP:conf/esws/VidalRLMSP10, DBLP:conf/www/StockerSBKR08, DBLP:conf/edbt/Gubichev014, DBLP:conf/kcap/LoustaunauH21, DBLP:conf/semweb/FerradaBH20].
@@ 7.2 Entailment Regimes and Query Processing
Although both approaches yield equivalent answers for a given SPARQL query, their performance can diverge significantly. Materialized RDF query processing may outperform on-the-fly execution of the rewritten query, but it may consume more memory [DBLP:conf/aaai/GlimmKKS15]. Nevertheless, various optimization techniques have been proposed to mitigate the overhead caused by the on-the-fly evaluation of entailment regimes [DBLP:journals/ws/TrivelaSCS15]. These optimizations are required particularly in the presence of the owl:sameAs. This predicate corresponds to logical equivalence and involves the application of the Leibniz Inference Rule [DBLP:books/sp/GriesS93] to deduce all the equivalent triples entailed by equivalent resources based on owl:sameAs relation. This process may lead to many intermediate results, impacting the query engine’s performance. Xiao et al. [DBLP:conf/esws/XiaoHBRGC18] propose query rewriting techniques to efficiently evaluate SPARQL queries with owl:sameAs employing equivalent SQL queries.
@@ 7.3 Federated Query Processing
Federations provide another perspective on querying linked data over multiple sources. A federation of knowledge graphs shares common entities while potentially providing different perspectives on those entities. Each knowledge graph within the federation operates autonomously and can be accessed through various Web interfaces, such as SPARQL endpoints or Linked Data Fragments (LDFs) [DBLP:journals/ws/VerborghSHHVMHC16]. SPARQL endpoints offer users the ability to execute any SPARQL query against multiple SPARQL endpoints. In contrast, LDFs enable access to specific graph patterns, such as triple patterns [DBLP:conf/semweb/VerborghHMHVSCCMW14a] or star-shaped graph patterns [DBLP:journals/corr/abs-2002-09172], allowing retrieval of fragments from an RDF knowledge graph. A star-shaped subquery is a conjunction of triple patterns in a SPARQL query that share the same subject variable [DBLP:conf/esws/VidalRLMSP10]. An LDF client can submit requests to a server, which then delivers results based on a data shipping policy and partitions results into batches of specified page sizes. Query processing in a federation of graphs differs from querying a single source because it enables real-time data integration of graphs from multiple sources. For example, Figure 10 depicts a SPARQL query whose execution requires the evaluation of subqueries over three knowledge graphs: a Cancer Knowledge Graph (CKG) [DBLP:journals/semweb/AisoposJNPRSIVM23], DBpedia, and Wikidata. This query could not be executed over a single data source unless the three knowledge graphs were physically materialized into one. Subqueries with a specific shape (e. g., star-shaped subqueries) need to be identified and posed against the knowledge graph(s) that is able to answer a particular part of the query. The federated query engine has to decompose input queries into these subqueries, find a plan to execute them and collect and merge the answers from the subqueries to produce a federated answer.
@@
A federated SPARQL query engine typically follows a mediator and wrapper architecture, which has been established in previous research [Wiederhold92, ZadorozhnyRVUB02]. Wrappers play a crucial role in translating SPARQL subqueries into requests sent to SPARQL endpoints, while also converting the endpoint responses into internal structures that the query engine can process. The mediator, on the other hand, is responsible for rewriting the original queries into subqueries that can be executed by the data sources within the federation. Additionally, the mediator collects and merges the results obtained from evaluating the subqueries to produce the final answer to the federated query. Essentially, the mediator consists of three main components:
@@
  • Source selection and query decomposition. This component decomposes queries into subqueries and selects the appropriate graphs (sources) capable of executing each subquery. Simple subqueries typically consist of a list of triple patterns that can be evaluated against at least one graph. Formally, source selection corresponds to the problem of finding the minimal number of knowledge graphs from the federation that can produce a complete answer to the input query. On the other hand, query decomposition requires partitioning the triple patterns of a query into a minimal number of subqueries, such that each subquery can be executed over at least one of the selected knowledge graphs. Commonly, federated query engines follow heuristic-based methods to solve these two problems. For example, for query decomposition, heuristics based on exclusive groups [DBLP:conf/semweb/SchwarteHHSS11] or star-shaped subqueries [DBLP:conf/esws/VidalRLMSP10, DBLP:journals/tlsdkcs/VidalCAMP16, DBLP:conf/semweb/MontoyaSH17] enable to efficiently solve source selection and query decomposition in queries free of general predicates (e. g., owl:sameAs or rdf:type).
@@
    Moreover, more general approaches (e. g., Endris et al. [DBLP:journals/tlsdkcs/EndrisGLMVA18]) resort to metadata describing the star-shaped patterns existing in a knowledge graph to perform query decomposition and source selection accurately. An exclusive group corresponds to a set of triple patterns (corresponding to a conjunction of the triples) in a SPARQL query that can be exclusively executed over one data source, as in FedX [DBLP:conf/semweb/SchwarteHHSS11]. Such a technique is implemented in the SPARQL engine FedX [DBLP:conf/semweb/SchwarteHHSS11], which also optimizes the execution of star-shaped subqueries by identifying groups of conjunctive triple patterns that a data source can execute independently from others.
@@
  • Query optimizer. This component identifies execution plans by combining star-shaped subqueries (SSQs) and utilizing physical operators implemented by the query engine. Formally, optimizing a query corresponds to the problem of finding a physical plan for the query that minimizes the values of a utility function (e. g., execution time or memory consumption). To maximize the utility function, query optimizers consider plans with different orders of executing operators, alternative implementations of operators, such as joins, as well as particular execution alternatives for certain query types, e. g., queries involving aggregation [DBLP:conf/esws/IbragimovHPZ15]. In general, finding an optimal solution is computationally intractable [DBLP:journals/tods/IbarakiK84], while the problems of constructing a bushy tree plan [DBLP:conf/pods/ScheufeleM97] and finding an optimal query decomposition over the graphs [DBLP:journals/tlsdkcs/VidalCAMP16] are NP-Hard. A bushy tree plan is a query execution plan that represents a query as a tree structure with multiple branches or subqueries, which can also be bushy-tree plans. Query plans can be generated following the traditional optimize-then-execute paradigm or re-optimize and adapt a plan on the fly according to the conditions and availability of selected graphs [DBLP:series/lncs/EndrisVG20]. Alternatively, the query optimizer may resort to a cost model to guide the search on the space of query plans and identify the one that minimizes the values of the utility function [DBLP:conf/semweb/MontoyaSH17].
@@
  • Query engine. This component of a federated query engine implements the physical operators necessary to combine tuples obtained from the graphs. These physical operators are designed to support logical SPARQL operations such as JOIN, UNION, or OPTIONAL [DBLP:journals/tods/PerezAG09]. Physical operators can be empowered to adapt execution schedulers to the current conditions of a group of selected graphs. Thus, adaptivity can be achieved at the intra-operator level, where the operators can detect when graphs become blocked or data traffic bursts. Additionally, intra-operator opportunistically produce results as quickly as data arrives from the graphs, and can produce results incrementally. Some opportunistic approaches [DBLP:conf/semweb/AcostaVLCR11, DBLP:conf/sigmod/HoseS12, DBLP:conf/i-semantics/GalkinEACVA17, DBLP:conf/semweb/AcostaV15] combine producing results quickly in an incremental fashion with greedy source selection so that the system stops querying additional graphs once the user’s wishes, e. g., in terms of the minimum number of obtained results, are fulfilled. The physical operators implemented by the SPARQL federated query engine ANAPSID [DBLP:conf/semweb/AcostaVLCR11] implement intra-operator adaptivity, enabling results generation even when the graphs became blocked. On the other hand, inter-operator adaptive strategies can produce an answer as soon as it is computed and can keep producing intermediate results even when data from a source becomes blocked. Acosta and Vidal [DBLP:conf/semweb/AcostaV15] propose adaptive query processing techniques that enable the re-ordering of a query execution plan to adjust execution schedulers to unexpected environmental conditions (e. g., changes at rates at which tuples arrive from graphs or a graph’s availability).
@@
The problem of query execution is defined in terms of data shipping strategies that are able to distribute query load between a server-client architecture of Linked Data Fragments. State of the art approaches tackle this issue by employing graph partitioning techniques to generate graph pattern fragments. These fragments aim to improve query performance, scalability, and workload balance. SAGE [DBLP:conf/f-ic/MinierSM19], smart-KG [DBLP:conf/www/AzzamFABP20], brTPF [DBLP:journals/corr/HartigA16], SkyTPF [DBLP:conf/semweb/KelesH19], SPF [DBLP:journals/corr/abs-2002-09172], TPF [DBLP:journals/ws/VerborghSHHVMHC16], and WiseKG [DBLP:conf/www/AzzamAMKPH21] are exemplary approaches for accessing graph patterns of varying shapes. These interfaces for Linked Data Fragments may significantly accelerate query execution, reduce workload, and minimize data transfer. Moreover, query rewriting based on these fragments correctly produces all the answers of the rewritten query, i. e., the methods are sound and complete. However, it is important to note that there exists a trade-off between memory consumption and execution time improvement so that storage requirements (e. g., smart-KG) can be twice as high in comparison to other approaches (e. g., SAGE, TPF, or SPF), due to the need to store partitions and their associated metadata.
Because of the variety of available interfaces, some works have proposed methods for heterogeneous federations [DBLP:conf/www/HelingA22, DBLP:conf/semweb/MontoyaKH19, DBLP:conf/semweb/MontoyaAH18a, DBLP:conf/esws/ChengH22], where the goal is to find efficient query execution plans in consideration of the constraints as well as strengths of the available interface.